Problem 1 XYZ Restaurant, only sells pepperoni pizzas. The restaurant incurs the following expenses: Each pizza sells for $10please answer one and two.
Answer #1

1) Fixed cost = 1500+3000+450 = 4950

Variable cost per pizza = 0.50+3+2 = 5.5

Contribution margin per pizza = 10-5.5 = 4.5 per pizza

2) Break even = Fixed cost/Contribution margin per pizza = 4950/4.5 = 1100 pizza

3) Required sales = (4950+6300)/4.5 = 2500 pizza

Note: Pls post each que individually as per HOMEWORKLIB RULES